# LATex: Leveraging Attribute-based Text Knowledge for Aerial-Ground Person Re-Identification

### v1

@@ 5.5 Ablation Studies
Figure 4: Performance with different number of prompts.
Figure 5: Accuracy of attribute predictions.
Figure 6: Rank list comparison of different models defined in Tab. 5. Correctly retrieved images are marked with a green box, while incorrect ones with a red box.
@@
Figure 8: Visualization of feature distributions with t-SNE(Van der Maaten & Hinton, 2008). Different colors refer to different attributes, each comprising several fine-grained subcategories perceptible by the corresponding PAC.
@@
Accuracy of Attribute Predictions. Attribute predictions play an important role in our framework. Fig. 5 shows the accuracy of some typical attributes predicted by PACG. It can be observed that our PACG achieves outstanding performances in these person attributes. These attributes provide discriminative information for person identification.
5.6 Visualization Analysis
Rank list Comparison. Fig. 6 compares the rank lists generated by different models defined in Tab. 5. With the sequential addition of PACG and CPT, the rank lists become progressively more accurate and discriminative. This indicates that our model gradually acquires the ability to perceive person attributes, enabling it to better distinguish individuals with similar visual characteristics.
@@ 5.6 Visualization Analysis
Attribute Retrieval. Fig. 7 illustrates the retrieval results using the attribute features of the query. Despite challenges such as image blurriness and small pixel regions of key areas, LATex accurately retrieves persons sharing a specific attribute (e.g., Upperwear) with the query’s ground truth. This result demonstrates the exceptional capability of LATex in perceiving person attributes.
@@
Attribute Feature Distributions. Fig. 8 shows the feature distribution of all attribute categories in the test set. Each attribute category consists of several subcategories, which are finely perceived by the corresponding PAC. For example, the “Hair style” category includes various styles such as “Bald”, “Short” and so on. The results clearly demonstrate that our method exhibits strong perception and discrimination capabilities for unseen samples across different attributes.
6 Conclusions
@@
In this paper, we propose a novel feature learning framework named LATex for AG-ReID. It adopts prompt-tuning strategies to leverage attribute-based text knowledge of visual-language models. To this end, we first propose an AIE to extract global semantic features and attribute-aware features. Then, we propose a PACG to generate person attribute predictions and obtain embedded tokens of predicted attributes. Finally, we design a CPT to transform attribute tokens and view information into structured sentences for more discriminative features. Extensive experiments on three AG-ReID benchmarks demonstrate superior performance of our methods. However, the experimental results still show a considerable gap compared to the ideal results presented in Fig. 2. We believe this may stem from insufficient accuracy in attribute predictions. In the future, improving person attribute perception with a higher precision could further advance research in AG-ReID.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendices
We organize the appendix as follows:
  • Appendix A : Details of the idealized assumption shown in Section 3.2 of the main paper.
  • Appendix B : Additional implementation details of the proposed model.
  • Appendix C : Extended experimental results, including numerical results and visualizations.
Appendix A Idealized Assumption Details
As elaborated in Section 3.2, the primary goal of the idealized assumption is to validate the feasibility of utilizing person attribute information as auxiliary data for cross-view person retrieval, as well as to explore the theoretical upper bound under a CLIP-based prompt-tuning framework. The network archtecture of idealized assumption is shown in Fig. 9(a). For the input images $\mathcal{C}^{Img}$, we consider prompt-tuning CLIP visual encoder, which provides the visual features $F_{v}$. For the textual input, we encode the ground truth of person attributes into a predefined template, specifically, “A photo of a \$Attribute\$ person,” to construct a contextualized text sequence. This sequence is then passed through the frozen CLIP text encoder to extract the corresponding textual feature $F_{t}$. $F_{v}$ and $F_{t}$ are supervised using the ReID loss defined in Eq. 11, respectively. During inference, the two features are concatenated as $[F_{t},F_{v}]$ to form a unified feature vector for retrieval.
As shown in Fig. 9(b), under the same backbone configuration as our proposed LATex, the idealized assumption achieves near-perfect performance on the AG-ReID.v1 dataset. These remarkable idealized experimental results motivate us to explore more practical approaches aligned with real-world scenarios.
(a) Overview of idealized assumption
(b) Experiment results of idealized assumption
Figure 9: Overview and experiment results of idealized assumption.
Appendix B More Implementation Details
We provide more details information of experiment implementation in this section:
LATex. In addition to the implementation details described in the main text, we utilize the Adam optimizer with a base learning rate of 0.00035 and a weight decay of 0.1 to optimize our model. A learning rate scheduling strategy is employed, combining a warm-up phase with cosine decay and a scaling factor of 0.01. Finally, in the AIE module, we leverage a larger number of learnable prompts for large-scale datasets to achieve better performance. Specifically, we use 128 prompts for AG-ReID.v2 and 256 for CARGO. Notably, despite the increased number of learnable prompts, only the first $T$ prompts, corresponding to the total number of person attributes, are used as attribute-aware prompts in subsequent training as described in the main text. The remaining prompts are solely utilized for fine-tuning the pre-trained model.
LATex†. LATex† is a variant of LATex that adopts the full fine-tuning strategy to ensure a fair comparison with other AG-ReID methods employing the same training strategy. Specifically, we unfroze all trainable parameters of the pre-trained CLIP vision and text encoders and updated them with a learning rate of 0.000005. Other experimental settings, including the learning rate decay strategy, optimizer, and the number of training epochs, are kept consistent with LATex.
Idealized Assumption. The idealized assumption represents a simplified version of LATex under ideal conditions, where only the backbone network of LATex is retained. All experimental settings are consistent with those of LATex.
Appendix C Extended Experiment
Table 8: Performance comparison evaluated using the mINP metric. The superscript symbol † indicates results based on full fine-tuning strategies. The best and second-best results are highlighted in bold and underline.
Method AG-ReID.v1 CARGO
[…]
ViT(Dosovitskiy et al., 2020) - - 39.62 28.20 57.55 47.07
VDT(Zhang et al., 2024b) 51.06 52.87 41.13 29.95 58.39 50.22
LATex 51.28 50.95 44.24 36.64 62.53 39.28
LATex† 54.27 56.50 54.06 45.96 68.45 54.77
Table 9: Performance evaluated using the mINP metric. The superscript symbol † indicates results based on full fine-tuning strategies.
Method AG-ReID.v2
A→C A→W C→A W→A
LATex 56.16 56.72 49.68 51.52
LATex† 63.43 61.87 55.84 56.74
C.1 More Performance Comparison
Tab. 8 compares the mINP performance of our model under various evaluation protocols on two AG-ReID benchmarks (AG-ReID.v1 and CARGO) with other Transformer-based methods. As the Explain (Nguyen et al., 2023b) neither reporte this metric in the paper nor release its code, we omit it from the table. For the same reason, Tab. 9 reports the mINP metric on AG-ReID.v2 exclusively for our proposed LATex and its variant LATex†. Consistent with the trends observed in the mAP and Rank-1 metrics discussed in the main text, LATex with a frozen backbone demonstrates superior aggregate performance. Meanwhile, with trainable backbone parameters, LATex† achieves the best results. Specifically, LATex† outperforms the previous best model, VDT, by over 13% on average across all protocols on the CARGO benchmark.
C.2 More Ablation Studies
Fig. 10 presents a more comprehensive ablation study on the effect of the number of learnable prompts on AG-ReID.v1 banchmark. Consistent with the conclusions we drawn in the main text, although slight fluctuations are observed, this factor is not a primary determinant of LATex’s performance.
(a) A→G
(b) G→A
Figure 10: Performance with different number of prompts on AG-ReID.v1.
C.3 More Visualization Analysis
Fig. 11 presents more detailed visualizations of retrieval results based on attribute features from query images. Despite challenges posed by complex visual conditions, LATex demonstrates robust fine-grained attribute perception performance.
Figure 11: Visualizations of retrieval results based on attribute features. Query images are highlighted with yellow boxes, while the corresponding attribute names and ground truth values are displayed in blue and green boxes, respectively.

### v2

@@ IV-D Training and Inference Cost Comparison
Figure 3: Performance with different numbers of prompts under two protocols.
Figure 4: Accuracy of attribute predictions in PACG.
IV-E Ablation Studies
To demonstrate the effectiveness of our proposed modules, we evaluate them on the AG-ReID.v1 dataset. The results are shown in Tab. V. Furthermore, we conduct comprehensive and evaluations to explore the details of our model design.
Effect of Different Modules. In Tab. V, Model A serves a our baseline. It achieves a Rank-1 of 81.69% and an mAP of 72.36% under the protocol A→G, while obtaining a Rank-1 of 83.89% and an mAP of 74.78% under the protocol G→A. With PACG, Model B increases the performance by 3.53% Rank-1 and 3.84% mAP under the protocol G→A. Notably, its performance is already better than previous methods, such as VDT. With CPT, Model C achieves the best result across all evaluation metrics. These results clearly demonstrate the effectiveness of our key modules.
Effect of Leveraging Person Attributes. We further validate the effectiveness of leveraging person attributes. As shown in Tab. VI, Model A and Model B are implemented without the CPT module. Model C and Model D are complete models. The key difference lies in the features used for retrieval. Specifically, $F_{a}$ denotes the concatenation of all $F_{a}^{t}$. As can be observed, the performance of Model B is superior to that of Model A. The performances of Model B and Model D are highly comparable. The performance of Model C is the best. These results clearly highlight the effectiveness of leveraging attribute-based text knowledge for AG-ReID.
Effect of Backbones. We emphasize that the core idea of our LATex is to extract person attributes, embed them as pseudo-texts, and feed them into a text encoder to achieve robust person ReID. With the strong vision-language alignment, we adopt CLIP as the backbone. Considering that our approach uses a different backbone from previous methods, we evaluate the effect of various backbones and report the performance on AG-ReID.v1 in Tab. VII. As can be seen, CLIP-based VDT and ViT-based LATex show a significant performance drop. The main reason is that they cannot fully leverage CLIP’s vision-language knowledge. These results clearly demonstrate that our LATex’s excellent performance stems from effectively utilizing CLIP’s vision-language knowledge, rather than its inherent encoding capabilities.
Effect of Trainable Prompts. Fig. 3 shows the effect of the number of learnable prompts $P^{i}_{v}$. With the increase of trainable parameters, the performance of LATex remains stable, demonstrating its robustness to this factor.
TABLE VIII: Performance with different numbers of shared tokens.
Number A→G G→A
Rank-1 mAP Rank-1 mAP
2 83.57 75.81 86.07 78.03
4 82.82 74.60 87.42 78.76
8 84.41 75.85 88.88 79.19
12 83.66 75.00 85.03 78.29
16 84.51 75.28 87.11 78.78
TABLE IX: Effectiveness of View Token. “VT” denotes the view token.
Method A→G G→A
Rank-1 mAP Rank-1 mAP
VDT 82.91 74.44 86.59 78.57
LATex(w/o VT) 83.94 75.13 87.11 78.30
LATex 84.41 75.85 88.88 79.19
Effect of Shared Tokens. Tab. VIII analyzes the impact of the number of shared tokens. We can observe that too few shared tokens would limit feature learning, while too many shared tokens may introduce noise. Based on the results of ablation studies, we set the number of shared tokens to 8.
Effect of View Token. Tab. IX shows the effectiveness of view token. Even though the variant without view token demonstrates competitive performance, explicitly incorporating view token helps LATex further enhance its cross-view retrieval capabilities. Since view token is camera-specific and can be pre-defined, we integrate this token into the CPT to achieve the optimal performance.
@@ IV-E Ablation Studies
Figure 6: Rank lists of different models defined in Tab. V. Correctly retrieved images are marked with a green box, while incorrect ones with a red box.
@@
Figure 7: Visualization of the attribute feature distributions with t-SNE[80]. Different colors refer to different attributes, each comprising several fine-grained subcategories.
@@
Accuracy Analysis of Attribute Predictions. Attribute predictions play an important role in our framework. Fig. 4 shows the accuracy of some typical attributes predicted by PACG. It can be observed that our PACG achieves outstanding performances in these person attributes. These attributes provide discriminative information for person ReID.
IV-F Visualization Analysis
@@
Attribute Query Retrieval. Fig. 5 illustrates the retrieval results using the attribute features as the query. Despite many challenges such as image blurriness and small key regions, LATex accurately retrieves persons sharing a specific attribute (e.g., upper clothes in our case). These results show the exceptional capability of LATex in perceiving person attributes.
Rank List Comparison. Fig. 6 compares the rank lists generated by different models, as defined in Tab. V. With the sequential addition of PACG and CPT, the rank lists become increasingly more accurate and discriminative. This indicates that our model progressively acquires the ability to perceive person attributes, enabling it to better distinguish individuals with similar visual characteristics.
@@
Attribute Feature Distributions. Fig. 7 illustrates the feature distributions of all attribute categories in the test set. Each attribute category consists of several subcategories, which are finely distinguished by the corresponding PAC. For example, the “Hair Style” category includes various subtypes, such as “Bald”, “Short”, and “Long”. The results provide evidence that our method exhibits strong perception and discrimination capabilities for unseen samples across diverse attributes.
V Conclusions
@@
In this paper, we propose a novel feature learning framework, named LATex, for AG-ReID. It adopts prompt-tuning strategies to integrate attribute-guided textual features from vision-language models. To this end, we first propose an AIE to extract global semantic features and attribute-aware features. Then, we propose a PACG to generate person attribute predictions and obtain representations of predicted attributes. Finally, we design a CPT to transform attribute representations and view information into structured sentences for more discriminative features. Extensive experiments on three AG-ReID benchmarks demonstrate the superior performance of our methods. In the future work, we will improve person attribute prediction to further advance research in AG-ReID.